he weak acid HQ has a pKa of 4.89. Calculate the [H3O+ ] of 0.020 M HQ.

1
Expert's answer
2020-11-02T09:05:57-0500

"pKa =-\log(Ka)\\
\log(Ka) = -pKa\\
Ka=10^{-pKa}\\
Ka= 10^{-4.89}\\
Ka = 1.288×10^{-5}"


"\begin{alignedat}{4}
&HQ_{(aq)}\quad + &\quad H_2O_{(l)}\quad _\longrightarrow^\longleftarrow\quad &H_3O^+_{(aq)}\quad + &\quad Q^-_{(aq)}
\end{alignedat}"

"\begin{aligned}
&\textsf{Initial Conc. : }0.020M +\ \text{---}M\ &= 0M + 0M\\
&\textsf{Eqm: }(0.020-x)M +\ \text{---}M\ &= xM + xM

\end{aligned}"


"\begin{aligned}
Ka = \dfrac{products}{reactants} = \dfrac{[H_3O^+][Q^-]}{HQ}
\end{aligned}"

"\begin{aligned}
1.288 × 10^{-5} = \dfrac{(x)(x)}{(0.020-x)}
\end{aligned}"


Since x < < 1,

0.020 - x "\approx" 0.020


"\begin{aligned}
1.288 × 10^{-5} = \dfrac{x^2}{(0.020)}
\end{aligned}"


"\begin{aligned}
x^2 = 1.288 × 10^{-5} × 0.020
\end{aligned}"


"\begin{aligned}
x = \sqrt{1.288 × 10^{-5} × 0.020\ }
\end{aligned}"

"x = 1.605 \u00d7 10^{-3} M"


"\begin{aligned}
[H_3O^+] = x = 1.605 × 10^{-3}M
\end{aligned}"